# Optimizing Parameters for Static Equilibrium of Discrete Elastic Rods with Active-Set Cholesky

### v1

@@ IV-D Newton-based Box-Constrained Optimization
$F(\Delta\mathbf{p})$ $=\frac{1}{2}\Delta\mathbf{p}^{T}\nabla^{2}L(\mathbf{p}^{k})\Delta\mathbf{p}+\Delta\mathbf{p}^{T}\nabla L(\mathbf{p}^{k},\bm{\lambda}^{k}),$ (21)
where $\Delta\mathbf{p}_{\mathrm{min}}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}$ denote lower and upper bounds for $\Delta\mathbf{p}$, respectively. We omit $\bm{\lambda}^{k}$ in the argument of $\nabla^{2}L(\mathbf{p})$ as it is independent of $\bm{\lambda}^{k}$ (23). Assuming the ideal step length of one [15], we have $\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}^{k+1}=\mathbf{p}^{k}+\Delta\mathbf{p}^{k+1}\leq\mathbf{p}_{\mathrm{max}}$, giving $\Delta\mathbf{p}_{\mathrm{min}}^{k}=\mathbf{p}_{\mathrm{min}}-\mathbf{p}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}=\mathbf{p}_{\mathrm{max}}-\mathbf{p}^{k}$ [42].
@@
$\nabla L(\mathbf{p},\bm{\lambda}^{k})=\nabla R(\mathbf{p})+\mathbf{J}^{T}\mathbf{M}^{-\frac{1}{2}}\bm{\lambda}^{k}+\rho\mathbf{J}^{T}\mathbf{M}^{-1}\mathbf{f}(\mathbf{p}),$ (22)
where $\nabla R(\mathbf{p})=\mathbf{W}(\mathbf{p}-\mathbf{p}_{0})$, and we denote the Jacobian of $\mathbf{f}(\mathbf{p})$ with respect to $\mathbf{p}$ by $\mathbf{J}\left(=\frac{\partial\mathbf{f}(\mathbf{p})}{\partial\mathbf{p}}\right)\in\mathbb{R}^{(4N-8)\times(7N-14)}$ (details are provided in Sec. IV-E). Note that we treat $\mathbf{M}$ (which is computed with the initial rest length $\mathbf{\bar{l}}$ [7, 40]) as constant, as $\mathbf{M}$ is introduced to properly scale forces (14).
Because the force formulations are primarily linear with respect to the rest shape and stiffness parameters, the second order derivatives of the forces with respect to $\mathbf{p}$ are mostly zero. An exception is the second order derivatives with respect to $\mathbf{\bar{l}}$; however, these components would be projected to zero to ensure SPD inner systems and valid Newton descent directions [15]. Thus, using the exact Hessian with SPD projections would have only negligible benefits while increasing the implementation complexity and computational cost due to the involvement of third order tensors [5]. Therefore, we ignore the second order derivatives with respect to $\mathbf{\bar{l}}$ and approximate the Hessian in a Gauss-Newton style (excluding $\bm{\lambda}$ from arguments):
@@
As this approximated Hessian is guaranteed to be SPD, we can solve the convex BCQP (20) using MPRGP [16], accelerated with our active-set Cholesky preconditioner (see Sec. V).
IV-E Forces and Jacobians
IV-E1 Inertia
Given the inertia force defined as $\mathbf{f}_{\mathrm{in}}=-\nabla E_{\mathrm{in}}(\mathbf{q})$ with $\mathbf{q}=\mathbf{q}^{t}$ and $\mathbf{\dot{q}}^{t}=0$ for a static equilibrium case, we have $\mathbf{f}_{\mathrm{in}}=\mathbf{f}_{\mathrm{ext}}$ and thus $\frac{\partial\mathbf{f}_{\mathrm{in}}}{\partial\mathbf{p}}=0$ [5].
IV-E2 Stretching
We define the stretching force of edge $i$ on vertex $i+1$ as $\mathbf{f}_{\mathrm{st,i,i+1}}=-\nabla_{\mathbf{x}_{i+1}}E_{\mathrm{st,i}}(\mathbf{x}_{i},\mathbf{x}_{i+1})$ (4), and define $\mathbf{f}_{\mathrm{st,i,i}}$ analogously. Given the dependence of $\mathbf{f}_{\mathrm{st,i,i+1}}$ and $\mathbf{f}_{\mathrm{st,i,i}}$ on $\mathbf{\bar{l}}_{i}$ and $\bm{\alpha}_{i}$, their Jacobians are
$\frac{\partial\mathbf{f}_{\mathrm{st,i,i+1}}}{\partial\mathbf{\bar{l}}_{i}}$ $=-\frac{\partial\mathbf{f}_{\mathrm{st,i,i}}}{\partial\mathbf{\bar{l}}_{i}}=s\bm{\alpha}_{i}\pi r^{2}\mathbf{l}_{i}\mathbf{\bar{l}}_{i}^{-2}\mathbf{t}_{i},$ (24)
$\frac{\partial\mathbf{f}_{\mathrm{st,i,i+1}}}{\partial\bm{\alpha}_{i}}$ $=-\frac{\partial\mathbf{f}_{\mathrm{st,i,i}}}{\partial\bm{\alpha}_{i}}=-s\pi r^{2}\left(\mathbf{l}_{i}\mathbf{\bar{l}}_{i}^{-1}-1\right)\mathbf{t}_{i}.$ (25)
IV-E3 Bending
We define the bending force according to (6) by $\mathbf{f}_{\mathrm{be},i}=-\nabla E_{\mathrm{be},i}(\mathbf{y}_{i})$. Given its dependence on $\mathbf{\bar{l}}_{i-1},\mathbf{\bar{l}}_{i}$, $\bm{\bar{\kappa}}_{i,0},\bm{\bar{\kappa}}_{i,1}$, and $\bm{\beta}_{i}$ while respecting the constraints $\bm{\bar{\kappa}}_{i,0}=\bm{\bar{\kappa}}_{i,2}$ and $\bm{\bar{\kappa}}_{i,1}=\bm{\bar{\kappa}}_{i,3}$, the Jacobians are given (with $j\in\{0,1\})$ by
$\frac{\partial\mathbf{f}_{\mathrm{be},i}}{\partial\mathbf{\bar{l}}_{i-1}}$ $=\frac{\partial\mathbf{f}_{\mathrm{be},i}}{\partial\mathbf{\bar{l}}_{i}}=\left(\frac{s\bm{\beta}_{i}\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})^{2}}\right)\mathbf{J}_{\mathrm{cu},i}^{T}(\bm{\kappa}_{i}-\bm{\bar{\kappa}}_{i}),$ (26)
$\frac{\partial\mathbf{f}_{\mathrm{be},i}}{\partial\bm{\bar{\kappa}}_{i,j}}$ $=\left(\frac{s\bm{\beta}_{i}\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\nabla\bm{\kappa}_{i,j}+\nabla\bm{\kappa}_{i,j+2}),$ (27)
$\frac{\partial\mathbf{f}_{\mathrm{be},i}}{\partial\bm{\beta}_{i}}$ $=-\left(\frac{s\pi r^{4}}{4(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)\mathbf{J}_{\mathrm{cu},i}^{T}(\bm{\kappa}_{i}-\bm{\bar{\kappa}}_{i}).$ (28)
IV-E4 Twisting
We define the twisting force as $\mathbf{f}_{\mathrm{tw},i}=-\nabla E_{\mathrm{tw},i}(\mathbf{y}_{i})$ (8). Given its dependence on $\mathbf{\bar{l}}_{i-1},\mathbf{\bar{l}}_{i}$, $\mathbf{\bar{m}}_{i}$, and $\bm{\gamma}_{i}$, the Jacobians are given by
$\frac{\partial\mathbf{f}_{\mathrm{tw},i}}{\partial\mathbf{\bar{l}}_{i-1}}$ $=\frac{\partial\mathbf{f}_{\mathrm{tw},i}}{\partial\mathbf{\bar{l}}_{i}}=\left(\frac{s\bm{\gamma}_{i}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})^{2}}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})\nabla\mathbf{m}_{i},$ (29)
$\frac{\partial\mathbf{f}_{\mathrm{tw},i}}{\partial\mathbf{\bar{m}}_{i}}$ $=\left(\frac{s\bm{\gamma}_{i}\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)\nabla\mathbf{m}_{i},$ (30)
$\frac{\partial\mathbf{f}_{\mathrm{tw},i}}{\partial{\bm{\gamma}}_{i}}$ $=-\left(\frac{s\pi r^{4}}{(\mathbf{\bar{l}}_{i-1}+\mathbf{\bar{l}}_{i})}\right)(\mathbf{m}_{i}-\mathbf{\bar{m}}_{i})\nabla\mathbf{m}_{i}.$ (31)
IV-F Algorithm and Implementation
Algorithm 1 outlines our parameter optimization method. To accelerate convergence and reduce the risk of converging to undesirable local minima, we use warm starting. We set $\rho=10^{6}$, and use $10^{3}$ and $10^{0}$ for the stiffness and rest shape parts of diagonals in $\mathbf{W}$, respectively, so that we prioritize achieving (in order) zero net forces, small changes in the stiffness parameters, and then rest shape parameters. We perform a single Newton iteration to solve our primal BCQP subproblem (12) as it serves as an inner problem within (11). To guarantee a decrease in the objective, we implement a backtracking line search [15]. The iterations are terminated when $\left\lVert\Delta\mathbf{p}\right\rVert_{2}$ falls below a threshold $\epsilon_{p}\ (=10^{-8})$, iteration count exceeds $k_{\mathrm{max}}\ (=100)$, or backtracking line search fails.
@@ IV-F Algorithm and Implementation
1:  Initialize generalized positions $\mathbf{q}$, radius $r$, stiffness coefficients $\mathbf{c}_{\mathrm{st}}$, $\mathbf{c}_{\mathrm{be}}$, $\mathbf{c}_{\mathrm{tw}}$, length $\mathbf{l}$, rest length $\mathbf{\bar{l}}$, generalized mass matrix $\mathbf{M}$, unit tangent vector $\mathbf{t}$, reference and material frames, curvature $\bm{\kappa}$, rest curvature $\bm{\bar{\kappa}}$, twist $\mathbf{m}$, rest twist $\mathbf{\bar{m}}$, stiffness scaling $s$, and material stiffness parameters $\bm{\alpha}$, $\bm{\beta}$, $\bm{\gamma}$, penalty parameter $\rho$, weight matrix $\mathbf{W}$, Lagrange multiplier $\bm{\lambda}^{k}=0$, iteration index $k=0$
@@
5:   Compute $\nabla L(\mathbf{p}^{k},\bm{\lambda}^{k})$, $\nabla^{2}L(\mathbf{p}^{k})$ with (22) and (23)
6:   Solve $\nabla^{2}L(\mathbf{p}^{k})\Delta\mathbf{p}^{k+1}=-\nabla L(\mathbf{p}^{k},\bm{\lambda}^{k})$ with $\Delta\mathbf{p}_{\mathrm{min}}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}$
7:   Update to $\mathbf{p}^{k+1}$ using backtracking line search with (15) and synchronization for $\bm{\bar{\kappa}}$
@@
10:  while $\epsilon_{p}<\left\lVert\Delta\mathbf{p}^{k+1}\right\rVert_{2}$, $k<k_{\mathrm{max}}$, &\&& line search succeeds
V Box-Constrained Quadratic Program Solver
@@
Box constraints are a specific type of inequality constraint that can be handled efficiently through active-set methods, without relying on barriers or penalties [15]. Since our BCQPs arise as primal subproblems within the constrained nonlinear minimization (11), we opt to solve them with a prescribed level of accuracy suitable for inexact Newton-like methods. This approach enables us to conserve computational resources by avoiding unnecessary precision [15]. Instead of employing direct solvers with active sets (e.g., Lemke’s method or Dantzig’s simplex algorithm), we prefer iterative methods that permit early termination. Given the stiff yet SPD inner systems, we employ MPRGP [16] and propose an effective new preconditioner based on full Cholesky factorization and triangular solves with active sets, leveraging the advantages of the banded system structures. We refer to this scheme as active-set Cholesky (ASC) preconditioning. Notably, since the Cholesky-based approach precisely solves a linear system, Cholesky-preconditioned MPRGP (or CG) finds a solution in just one iteration when no box constraints are activated.
While Narain et al. [43] used modified incomplete Cholesky (MIC) preconditioning [44] for MPRGP, the derivations and rationale behind their approach are not explained, and it is not necessarily clear from their publicly available source code how to adapt it for variants of Cholesky decomposition (e.g., LDLT-type decomposition with $2\times 2$ diagonal blocks [45]). We therefore first discuss issues associated with the naive application of Cholesky factors and active sets, followed by an explanation of our methodology, including practical implementation details (Sec. V-A). We also clarify the differences between active-set IC preconditioning [43] and our scheme (Sec V-B).
V-A Active-Set Cholesky Preconditioning
@@
Consider the minimization of a quadratic objective $\frac{1}{2}\mathbf{x}^{T}\mathbf{A}\mathbf{x}-\mathbf{x}^{T}\mathbf{b}$ and corresponding linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$. The associated linear preconditioning system can be expressed as $\mathbf{A}\mathbf{z}=\mathbf{r}$, where $\mathbf{z}$ and $\mathbf{r}$ are vectors corresponding to $\mathbf{x}$ and $\mathbf{b}$, respectively. Adopting an active set $\mathbf{a}$ that indicates whether $\mathbf{x}$ is constrained or not (i.e., with an index $i$, $\mathbf{a}_{i}=1$ or $\mathbf{a}_{i}=-1$ if $\mathbf{x}_{i}$ is constrained by its lower or upper bound, respectively, and $\mathbf{a}_{i}=0$ otherwise) [46], we can define a diagonal selection matrix $\mathbf{S}$ with $\mathbf{S}_{ii}=1-|\mathbf{a}_{i}|$. Given $\mathbf{z}_{u}$ and $\mathbf{z}_{c}$ that denote unconstrained and constrained parts of $\mathbf{z}$, respectively, we need to solve the preconditioning system for $\mathbf{z}_{u}$ while enforcing $\mathbf{z}_{c}=0$ [46]. The preconditioning system can be rewritten (incorporating the necessary constraints [47]) as $(\mathbf{S}\mathbf{A}\mathbf{S}^{T}+\mathbf{I}-\mathbf{S})\mathbf{z}=\mathbf{S}\mathbf{r}$, or given equivalently (by splitting it for $\mathbf{z}_{u}$ and $\mathbf{z}_{c}$) as $\mathbf{S}_{u}\mathbf{A}\mathbf{S}_{u}^{T}\mathbf{z}_{u}=\mathbf{S}_{u}\mathbf{r}$ and $\mathbf{z}_{c}=0$, where $\mathbf{S}_{u}$ denotes a matrix consisting of $\mathbf{S}$’s rows associated with unconstrained variables. One approach to solving these preconditioning systems is to reassemble $(\mathbf{S}\mathbf{A}\mathbf{S}^{T}+\mathbf{I}-\mathbf{S})$ or $\mathbf{S}_{u}\mathbf{A}\mathbf{S}_{u}^{T}$ and factorize it using Cholesky decomposition, but this procedure is costly since active sets can change in each MPRGP iteration [46]. We would rather perform Cholesky decomposition only once and reuse its resulting factor.
V-A1 Problems with Naive Substitution of Cholesky Factors

### v2

@@ V-E Newton-based Box-Constrained Optimization
$F(\Delta\mathbf{p})$ $=\frac{1}{2}\Delta\mathbf{p}^{T}\nabla^{2}L(\mathbf{p}^{k})\Delta\mathbf{p}+\Delta\mathbf{p}^{T}\nabla L(\mathbf{p}^{k},\boldsymbol{\lambda}^{k}),$ (17)
where $\Delta\mathbf{p}_{\mathrm{min}}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}$ denote lower and upper bounds for $\Delta\mathbf{p}$, respectively. We omit $\boldsymbol{\lambda}^{k}$ in the argument of $\nabla^{2}L(\mathbf{p})$ as it is independent of $\boldsymbol{\lambda}^{k}$ (19). Assuming the ideal step length of one [15], we have $\mathbf{p}_{\mathrm{min}}\leq\mathbf{p}^{k+1}=\mathbf{p}^{k}+\Delta\mathbf{p}^{k+1}\leq\mathbf{p}_{\mathrm{max}}$, giving $\Delta\mathbf{p}_{\mathrm{min}}^{k}=\mathbf{p}_{\mathrm{min}}-\mathbf{p}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}=\mathbf{p}_{\mathrm{max}}-\mathbf{p}^{k}$ [50].
@@
$\nabla L(\mathbf{p},\boldsymbol{\lambda}^{k})=\nabla R(\mathbf{p})-\mathbf{J}^{T}\mathbf{M}^{-\frac{1}{2}}\boldsymbol{\lambda}^{k}+\rho\mathbf{J}^{T}\mathbf{M}^{-1}\mathbf{f}(\mathbf{p}),$ (18)
where $\nabla R(\mathbf{p})=\mathbf{W}(\mathbf{p}-\mathbf{p}_{0})$, and we denote the Jacobian of $\mathbf{f}(\mathbf{p})$ with respect to $\mathbf{p}$ by $\mathbf{J}\left(=\frac{\partial\mathbf{f}(\mathbf{p})}{\partial\mathbf{p}}\right)\in\mathbb{R}^{(4N-8)\times(7N-14)}$ (details are given in the supplementary material). Note that we treat $\mathbf{M}$ (which is computed with the initial rest length $\mathbf{\bar{l}}$ [7, 46]) as constant, as $\mathbf{M}$ is introduced and intended to properly scale the generalized forces (see (10)) [5].
Since the generalized forces are linear with respect to the rest curvature, rest twist, and stiffness parameters, the second order derivatives with respect to these parameters are zero. While the second order derivatives with respect to $\mathbf{\bar{l}}$ are non-zero, these components would be projected to zero to ensure SPD inner systems and valid Newton descent directions [15]. In addition, the change in $\mathbf{\bar{l}}$ during the parameter optimization is typically small for less extensible strands, i.e., $\mathbf{\bar{l}}$ can be close to constant. Thus, using the exact Hessian with SPD projections does not have discernible benefits while increasing the implementation complexity and computational cost due to the involvement of third order tensors, as demonstrated and discussed in [5]. Therefore, we ignore the second order derivatives with respect to $\mathbf{\bar{l}}$ and approximate the Hessian in a Gauss-Newton style (excluding $\boldsymbol{\lambda}$ from arguments) by
@@
As this approximated Hessian is guaranteed to be SPD, we can solve the convex BCQP (16) using MPRGP [16], accelerated with our active-set Cholesky preconditioner (see Sec. VI).
V-F Algorithm and Implementation
Algorithm 1 outlines our parameter optimization method. To accelerate convergence and reduce the risk of converging to undesirable local minima, we use warm starting. We set $\rho=10^{6}$, and use $10^{3}$ and $10^{0}$ for the stiffness and rest shape parts of diagonals in $\mathbf{W}$, respectively, so that we prioritize achieving (in order) zero net forces, small changes in the stiffness parameters, and then in rest shape parameters. We perform a single Newton iteration to solve our primal BCQP subproblem (8) as it serves as an inner problem within (7). To guarantee a decrease in the objective, we implement a backtracking line search [15]. The iterations are terminated when $\left\lVert\Delta\mathbf{p}\right\rVert_{2}$ falls below a threshold $\epsilon_{p}\ (=10^{-8})$ or iteration count exceeds $k_{\mathrm{max}}\ (=100)$. In addition, we can also terminate the solver iterations to prevent infinite loops if we cannot make any progress with the backtracking line search (although our method did not experience this case in our examples).
@@ V-F Algorithm and Implementation
1: Initialize generalized positions $\mathbf{q}$, radius $r$, stiffness coefficients $\mathbf{c}_{\mathrm{st}}$, $\mathbf{c}_{\mathrm{be}}$, $\mathbf{c}_{\mathrm{tw}}$, length $\mathbf{l}$, rest length $\mathbf{\bar{l}}$, generalized mass matrix $\mathbf{M}$, unit tangent vector $\mathbf{t}$, reference and material frames, curvature $\boldsymbol{\kappa}$, rest curvature $\boldsymbol{\bar{\kappa}}$, twist $\mathbf{m}$, rest twist $\mathbf{\bar{m}}$, stiffness scaling $s$, and material stiffness parameters $\boldsymbol{\alpha}$, $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$, penalty parameter $\rho$, weight matrix $\mathbf{W}$, Lagrange multiplier $\boldsymbol{\lambda}^{k}=0$, iteration index $k=0$
@@
5:  Compute $\nabla L(\mathbf{p}^{k},\boldsymbol{\lambda}^{k})$, $\nabla^{2}L(\mathbf{p}^{k})$ with (18) and (19)
6:  Solve $\nabla^{2}L(\mathbf{p}^{k})\Delta\mathbf{p}^{k+1}=-\nabla L(\mathbf{p}^{k},\boldsymbol{\lambda}^{k})$ with $\Delta\mathbf{p}_{\mathrm{min}}^{k}$ and $\Delta\mathbf{p}_{\mathrm{max}}^{k}$
7:  Update to $\mathbf{p}^{k+1}$ using backtracking line search with (11) and synchronization for $\boldsymbol{\bar{\kappa}}$
@@
10:while $\epsilon_{p}<\left\lVert\Delta\mathbf{p}^{k+1}\right\rVert_{2}$ and $k<k_{\mathrm{max}}$
VI Box-Constrained Quadratic Program Solver
@@
Box constraints are a specific type of inequality constraint that can be handled efficiently through active-set methods, without relying on barriers or penalties [15]. Since our BCQPs arise as primal subproblems (8) within the constrained nonlinear minimization (7), we opt to solve them with a prescribed level of accuracy suitable for inexact Newton-like methods. This approach enables us to conserve computational resources by avoiding unnecessary precision [15]. Thus, instead of employing direct solvers with active sets (e.g., Lemke’s method or Dantzig’s simplex algorithm), we prefer iterative methods that permit early termination. Given the stiff yet SPD inner systems, we employ MPRGP [16] and propose an effective new preconditioner based on full Cholesky factorization and triangular solves with active sets, utilizing the banded system structures. We refer to this scheme as active-set Cholesky (ASC) preconditioning. Notably, since the Cholesky-based approach precisely solves a linear system, Cholesky-preconditioned MPRGP (or CG) finds a solution in just one iteration when no box constraints are activated.
VI-A Active-Set Cholesky Preconditioning
@@
Consider minimizing a quadratic objective $\frac{1}{2}\mathbf{x}^{T}\mathbf{A}\mathbf{x}-\mathbf{x}^{T}\mathbf{b}$ and corresponding linear system $\mathbf{A}\mathbf{x}=\mathbf{b}$. The associated linear preconditioning system can be expressed as $\mathbf{A}\mathbf{z}=\mathbf{r}$, where $\mathbf{z}$ and $\mathbf{r}$ are vectors corresponding to $\mathbf{x}$ and $\mathbf{b}$, respectively. Adopting an active set $\mathbf{a}$ that indicates whether $\mathbf{x}$ is limited or not (i.e., with an index $i$, $\mathbf{a}_{i}=1$ or $\mathbf{a}_{i}=-1$ if $\mathbf{x}_{i}$ is limited by its lower or upper bound, respectively, and $\mathbf{a}_{i}=0$ otherwise) [44], we can define a diagonal selection matrix $\mathbf{S}$ with $\mathbf{S}_{ii}=1-|\mathbf{a}_{i}|$, which can also be interpreted as a filtering matrix. Given $\mathbf{z}_{u}$ and $\mathbf{z}_{c}$ that denote unconstrained and constrained parts of $\mathbf{z}$, respectively, we need to solve the preconditioning system for $\mathbf{z}_{u}$ while enforcing $\mathbf{z}_{c}=0$ [44]. The preconditioning system can be rewritten (incorporating the necessary constraints [51]) as $(\mathbf{S}\mathbf{A}\mathbf{S}^{T}+\mathbf{I}-\mathbf{S})\mathbf{z}=\mathbf{S}\mathbf{r}$, or given equivalently (by splitting it for $\mathbf{z}_{u}$ and $\mathbf{z}_{c}$) as $\mathbf{S}_{u}\mathbf{A}\mathbf{S}_{u}^{T}\mathbf{z}_{u}=\mathbf{S}_{u}\mathbf{r}$ and $\mathbf{z}_{c}=0$, where $\mathbf{S}_{u}$ denotes a matrix consisting of $\mathbf{S}$’s rows associated with the unconstrained variables.
VI-A1 Problems with Naive Application of Cholesky Solver
While one approach to solving these preconditioning systems is to reassemble $(\mathbf{S}\mathbf{A}\mathbf{S}^{T}+\mathbf{I}-\mathbf{S})$ or $\mathbf{S}_{u}\mathbf{A}\mathbf{S}_{u}^{T}$ and factorize it using Cholesky decomposition, this procedure is costly because active sets can change in each MPRGP iteration [44], and thus it is ideal to perform Cholesky decomposition only once and reuse its resulting factor.